Suppose you just received a shipment of twelve televisions. Three of the televisions are defective. If two televisions are randomly​ selected, compute the probability that both televisions work. What is the probability at least one of the two televisions does not​ work?
1
Expert's answer
2019-07-12T09:33:24-0400

the probability that the first selected TV is working 9/12=3/4

the probability that the second selected TV is working (9-1)/(12-1)=8/11

probability that both televisions work (3/4)*(8/11)=24/44=6/11

probability at least one of the two televisions does not​ work 1-(6/11)=5/11